OpenStudy (anonymous):

what is the equation of a line that has a slope of -3/2 and passes through the point (-4,2)?

OpenStudy (anonymous):

\[y=-3/2x-4\]

OpenStudy (anonymous):

you have to use slope intercept form y=mx+b m is slope so you just put in -3/2 and b is the y value that you have which is 2 so -3/2x+2

OpenStudy (anonymous):

b is not just the y value that you are given. You have to substitute in the x into the equation so you get 2=-3/2(-4)+b which gives you b=-4. Then you plug that in as your b value

OpenStudy (anonymous):

Thank you! (:

OpenStudy (anonymous):

You are welcome

OpenStudy (anonymous):

oh